A 7.65 kg bowling ball traveling at 12.65 m/s strikes a 1.7 kg bowling pin. The impact slows the bowling ball to 10.13 m/s.



Determine the change in momentum of the pin.
1
Expert's answer
2020-02-03T14:00:42-0500
"P=M(v_f-v_i)"

"P=7.65(12.65-10.13)=19.3\frac{kg\cdot m}{s}"